Metal sphere A has a charge of -Q. An identical me
Answer #1

force on B due to A=k*charge on B*charge on A/(distance between A and B)^2

where k=coloumb's constant=9*10^9


hence force on B due to A=k*2*Q*(-Q)/AB^2=-2*k*Q^2/AB^2

force on A due to B=k*charge on A*charge on B/(distance between A and B)^2=k*2*Q*(-Q)/AB^2=-2*k*Q^2/AB^2

hence both the forces have same magnitude.

so option 4 is correct.

force magnitude=F